OpenStudy (anonymous):

Math Analysis: Sketch the polar function 16. r = 1-cos 0

OpenStudy (anonymous):

It is a cardioid. Why?

OpenStudy (anonymous):

idk what a cardioid is

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=PolarPlot [+1+-+Cos[\[Theta]]%2C+{\[Theta]%2C+0%2C+2+\[Pi]}]

OpenStudy (anonymous):

Copy and past the whole link above in your browser

OpenStudy (anonymous):

ohh okay :) but i just need to know how to plot it for my test

OpenStudy (anonymous):

and show my work

OpenStudy (anonymous):

Take several points for different value of \[ \theta\] \[ \theta=0,\, r =0\\ \theta=\pm \frac \pi 3,\, r =\frac 1 2\\ \] etc and plot these points in polar coordinates.

OpenStudy (anonymous):

this is what my teacher did \[\theta | 0 | 45| 90| 135|180|225|270|315|360 \ \[r|0|0.3|1|1.7|2|1.7|1|0.3|0\]

OpenStudy (anonymous):



OpenStudy (anonymous):

he put that info on the chart i can't really draw it in sorry

OpenStudy (anonymous):

See the cardiode with the points for \[ \theta \in \left\{0,\frac{\pi }{6},\frac{\pi }{3},\frac{\pi }{2},\frac{2 \pi }{3},\frac{5 \pi }{6},\pi ,\frac{7 \pi }{6},\frac{4 \pi }{3},\frac{3 \pi }{2},\frac{5 \pi }{3},\frac{11 \pi }{6},2 \pi \right\} \]

OpenStudy (anonymous):

Did you get it now?

OpenStudy (anonymous):

how did you get those points thow

OpenStudy (anonymous):

my teacher did it by degrees

OpenStudy (anonymous):

would i just continually add 45 degrees up until 360 to plot my points?

OpenStudy (anonymous):

If you want more points do it every 30 degrees, like I did for my graph above. Where you able to see the graph.?

OpenStudy (anonymous):

wait so how did you automatically know that the equation was a cardioid

OpenStudy (anonymous):

my teacher did every 45 degrees

OpenStudy (anonymous):

That is fine, you get fewer points. \[ r = 1 \pm \sin(\theta)\\ r = 1 \pm \cos(\theta)\\ \] are cardioid

OpenStudy (anonymous):

cardioids look like those bean like graphs rite?

OpenStudy (anonymous):

cardioids look like heart. Have you heard about cardiac arrest?

OpenStudy (anonymous):

yes i have lol

OpenStudy (anonymous):

ohh i get it now :)

OpenStudy (anonymous):

so then, since I know that the equation is a cardioid it would look like a heart but then how did my teacher know how to plot the points for it?

OpenStudy (anonymous):

If you plot several points you get the shape of the heart. Try ploting \[ r = 1 + \sin(\theta) \]

OpenStudy (anonymous):

so i just randomly guess my points..? o-o

OpenStudy (anonymous):

by sketching the graph

OpenStudy (anonymous):

oh sorry :( i will gives you in a minute ;33

OpenStudy (anonymous):

You take some values of theta that you know the cosine of and you compute the corresponding r and you plot them.

OpenStudy (anonymous):

so if i want my theta values to be every 45 degrees i would do what again..?

OpenStudy (anonymous):

sorry \[ (\theta, r) \in \left( \begin{array}{cc} 0 & 0 \\ \frac{\pi }{4} & 1-\frac{1}{\sqrt{2}} \\ \frac{\pi }{2} & 1 \\ \frac{3 \pi }{4} & 1+\frac{1}{\sqrt{2}} \\ \pi & 2 \\ \frac{5 \pi }{4} & 1+\frac{1}{\sqrt{2}} \\ \frac{3 \pi }{2} & 1 \\ \frac{7 \pi }{4} & 1-\frac{1}{\sqrt{2}} \\ 2 \pi & 0 \\ \end{array} \right) \]

OpenStudy (anonymous):

what if i get an equation and I can't guess what it is right away?

OpenStudy (anonymous):

and i can't guess what shape of graph it is

OpenStudy (anonymous):

where did u get 1/square root of 2? o-o

OpenStudy (anonymous):

Take more points to try to guess the sahpe.

OpenStudy (anonymous):

\[ \cos(45) = \frac 1 {\sqrt 2} = \frac {\sqrt 2} 2 \]

OpenStudy (anonymous):

why don't you just leave it at square root of 2/2? o:

OpenStudy (anonymous):

would i get the cosine of the degree or the sine?

OpenStudy (anonymous):

It is the same number.

OpenStudy (anonymous):

cosine and sine are equal when the angle is 45

OpenStudy (anonymous):

how about when its 135 degrees

OpenStudy (anonymous):

cos( 135) =cos(45)

OpenStudy (anonymous):

Try it with your calculator

OpenStudy (anonymous):

Have you had trigonometry before?

OpenStudy (anonymous):

yeah. im in it atm actually lol

OpenStudy (anonymous):

ok thx i get it now! :P